OpenStudy (anonymous):

Find the derivative of -(1)/(20000)x^2

OpenStudy (sleepyhead314):

you can rewrite it as -(20000x^2)^(-1) then do power rule + chain so derivative = -(-1)(20000x^2)^(-2) * (20000(2x)) then simplify

OpenStudy (anonymous):

Is the answer -0.0001?

OpenStudy (sleepyhead314):

??? I think there is still supposed to be an x in there

OpenStudy (anonymous):

opps forgot the x

OpenStudy (sleepyhead314):

negatives cancel out?

OpenStudy (sleepyhead314):

and it is 0.0001 (x/x^4) = 0.0001(x^-3)

OpenStudy (anonymous):

ahh ok I see now.

OpenStudy (sleepyhead314):

oh wait... does the x start on the top or bottom?

OpenStudy (anonymous):

\[-\frac{ 1 }{ 20000 }x^2\]

OpenStudy (sleepyhead314):

sorry, I had did it as on the bottom here is the correct wolfram alpha answer http://www.wolframalpha.com/input/?i=derivative+of+-%281%29%2F%2820000%29x%5E2+

OpenStudy (anonymous):

ahh ok in the computer they have it written as -0.0001x I was wondering how did they get that. from \[-\frac{ x }{ 10000 }\]

OpenStudy (anonymous):

nvm got it now hehe thanks